# Pruning Large Language Models with Semi-Structural Adaptive Sparse Training

### v2

@@ 4.1 Experiment Setup
Model Configurations. We report the performance of Adaptive Sparse Trainer (AST) across three different LLM model families: LLaMA2-7B (Touvron et al. 2023), OPT (Zhang et al. 2022), and GPT2 (Brown et al. 2020). We present results for two different sparsity patterns: AST-Naive, which adheres to a strict 2:4 sparsity pattern without additional weights, and AST-Boosted, which incorporates extra SLoRB weights. For AST-Boosted models, we selected $k=16$, introducing an additional 12.5% of parameters. Optimal hyperparameters were identified through a grid search, with the specific hyperparameters and training details provided in Section 6.3.
Data. For training smaller models like the OPT and GPT2 model families, we utilized the C4 dataset (Raffel et al. 2020). For the LLaMA2-7B model, we employed a more comprehensive dataset, RedPajama-v1111https://huggingface.co/datasets/togethercomputer/RedPajama-Data-1T, which encompasses data from seven domains: CommonCrawl, C4, GitHub, Wikipedia, Books, ArXiv, and StackExchange. Additionally, we leveraged the dynamic batching feature provided in the ShearedLLaMA (Xia et al. 2024) codebase.
@@ 4.2 Main Results
Our method can also be seamlessly adapted to current quantization techniques to achieve extremely compressed models with minimal performance drop. We provide results using AWQ quantization methods in section 6.8.
@@ 4.4 Ablation Study
Distillation Functions We also conducted ablation studies on different distillation functions used in previous work, including TinyBERT (Jiao et al. 2020), MobileBERT (Sun et al. 2020), and Sparse-Finetuning (Kurtic et al. 2023), which use attention and hidden states for distillation. Additionally, we examined MiniLLM (Gu et al. 2024), which employs reverse KL divergence. We find that using intermediate information during the distillation of generative language models is detrimental, and that KL loss is sufficient for optimal performance. Reverse-KL yields performance similar to that of forward KL. Detailed descriptions of the distillation loss functions for each of these methods are provided in section 6.5.
Table 5: Ablation study of different distillation loss for training sparse models.
@@ 4.5 Speedup
In Table LABEL:tab6, we report speedup results using PyTorch support for 2:4 sparse models [^2] https://pytorch.org/docs/2.1/sparse.html#sparse-semi-structured-tensors. Since the current API only accelerates computation without reducing memory consumption, we focus on reporting results for matrix multiplication and end-to-end inference on LLaMA2-7B. 2:4 sprase model can achieves up to a 1.91x speedup in matrix multiplication and up to 1.35x in end-to-end inference.
Table 6: Speedup results using PyTorch API. We fixed the size of the sparse matrix to be 4096 by 4096 and varied the input matrix size when testing matrix multiplication.
Matrix Multiplication
Block Size Sparse Dense Speedup
12288 3.80ms 1.99ms 1.91x
8096 2.54ms 1.33ms 1.91x
4096 1.28ms 0.68ms 1.88x
End-to-End
Seqlen Sparse Dense Speedup
2048 289.07ms 244.80ms 1.18x
1024 142.04ms 105.22ms 1.35x
512 73.84ms 55.68ms 1.35x
@@ 5 Conclusion
In this paper, we introduce the Adaptive Sparse Trainer (AST), a novel training pipeline designed for semi-structured sparse models. AST effectively narrows the precision gap between dense and sparse large language models (LLMs), both in terms of perplexity and performance on knowledge-intensive tasks, while maintaining minimal training costs. Our work demonstrates the feasibility of pruning LLMs without significant performance loss, showing that large semi-structured sparse models can outperform dense models of similar size when paired with appropriate hardware. Additionally, AST achieves substantial performance gains without compromising inference speed, highlighting its potential for the efficient deployment of sparse LLMs. While our findings contribute to the advancement of retraining pruned models with billions of parameters, we acknowledge that the use of a limited number of training tokens due to computational constraints is an area for future exploration. Extending this work to larger models or increasing the number of training tokens could provide valuable insights and further enhance the effectiveness of AST.
@@ 6.1 Compressing Models with Different Semi-structured Sparsity Pattern
Through mathematical formulation (provided in Section 6.2), we can provide an approximate upper bound for the compression ratio as well as a predicted actual compression ratio when extending to sparsity patterns with larger $N$ and $M$
@@ 6.5 Different Distillation Function
6.6 Adding Decay for Adam Optimizer
Recent work (Hu et al. 2024) has discovered that using SR-STE with momentum-based optimizers can diminish the effectiveness of the decay term, leading to suboptimal results. Previously, the decay term was added to the gradient as follows:
$g_{t}\leftarrow\tilde{g_{t}}+\lambda_{W}(\overline{m(W_{t})}\odot W_{t}),$ (16)
$W_{t+1}\leftarrow W_{t}-\frac{\gamma(u_{t}\beta_{1}+(1-\beta_{1})g_{t})}{(1-\beta_{1}^{t})(\sqrt{\hat{v_{t}}}+\epsilon)},$ (17)
where $\tilde{g_{t}}$ is the gradient of $\tilde{W_{t}}$, and $u_{t}$ and $v_{t}$ are the first-order and second-order momenta, respectively.
Now, we add the decay term directly to the weight. First, we calculate the gradient $g(\tilde{W_{t}})$ using the Adam optimizer, and then update the weight using:
$W_{t+1}\leftarrow W_{t}-\gamma_{t}\left(g(\tilde{W_{t}})+\lambda_{W}(\overline{m(W_{t})}\odot W_{t})\right).$ (18)
@@ 6.7 Mask Oscillation Behavior During Training
Table 10: Wikitext perplexity for quantized sparse models.
LLAMA2 OPT GPT2
Method
Theoretical
Compression Rate
7B 125M 1.3B 124M 350M 774M 1.5B
Dense 1.0x 5.12 27.76 14.62 29.95 21.72 19.43 17.40
AWQ-4bit 0.125x 5.68 29.12 14.95 31.93 22.66 19.89 17.80
AWQ-2bit 0.0675x 2.2e5 216.11 53.42 751.15 453.45 70.36 46.17
AST-Naive-8bit
0.125x 6.37 30.26 15.86 32.36 23.66 21.29 18.34
AST-Naive-4bit
0.0675x 6.48 31.28 16.05 34.00 24.39 21.72 18.60
AST-Boosted-4bit
0.078x 6.25 29.88 15.63 32.76 23.87 21.40 18.31
@@ 6.8 Model Compression Using AWQ
In this section, we present the perplexity results for quantized models using AWQ. As shown in Table 10, our model maintains performance even under extreme compression. For the AST-Boosted model, we also quantized the additional SLoRB parameters.

### v3

@@ 4.1 Experiment Setup
Model Configurations. We report the performance of Adaptive Sparse Trainer (AST) across three different LLM model families: LLaMA2-7B (Touvron et al. 2023), OPT (Zhang et al. 2022), and GPT2 (Brown et al. 2020). We present results for two different sparsity patterns: AST-Naive, which adheres to a strict 2:4 sparsity pattern without additional weights, and AST-Boosted, which incorporates extra SLoRB weights. For AST-Boosted models, we selected $k=16$, introducing an additional 12.5% of parameters. Optimal hyperparameters were identified through a grid search, with the specific hyperparameters and training details provided in Section 3 of the Appendix.
Data. For training smaller models like the OPT and GPT2 model families, we utilized the C4 dataset (Raffel et al. 2020). For the LLaMA2-7B model, we employed a more comprehensive dataset, RedPajama-v1222https://huggingface.co/datasets/togethercomputer/RedPajama-Data-1T, which encompasses data from seven domains: CommonCrawl, C4, GitHub, Wikipedia, Books, ArXiv, and StackExchange. Additionally, we leveraged the dynamic batching feature provided in the ShearedLLaMA (Xia et al. 2024) codebase.
@@ 4.2 Main Results
Our method can also be seamlessly adapted to current quantization techniques to achieve extremely compressed models with minimal performance drop. We provide results using AWQ quantization methods in section 8 of the Appendix.
@@ 4.4 Ablation Study
Distillation Functions We also conducted ablation studies on different distillation functions used in previous work, including TinyBERT (Jiao et al. 2020), MobileBERT (Sun et al. 2020), and Sparse-Finetuning (Kurtic et al. 2023), which use attention and hidden states for distillation. Additionally, we examined MiniLLM (Gu et al. 2024), which employs reverse KL divergence. We find that using intermediate information during the distillation of generative language models is detrimental, and that KL loss is sufficient for optimal performance. Reverse-KL yields performance similar to that of forward KL. Detailed descriptions of the distillation loss functions for each of these methods are provided in Section 5 of the Appendix.
Table 5: Ablation study on different distillation loss for training sparse models.
@@ 4.5 Speedup
Table LABEL:tab6 presents speedup results obtained using TensorRT-LLM [^3] https://github.com/NVIDIA/TensorRT-LLM. We evaluate the actual end-to-end decoding speedup on two GPU architectures with the LLaMA2-7B 2:4 sparse model. We employ throughput, measured as the number of tokens processed per second, as the primary evaluation metric. Across a range of input and output lengths, the 2:4 sparse model demonstrates an overall acceleration of 1.33× to 1.83× compared to its dense counterpart, highlighting its potential for practical deployment.
Table 6: Speedup results using TensorRT-LLM on RTX4090 and L20 GPUs with different input and output sequence lengths, measured by throughput (tokens/s).
RTX4090
Inp Len, Out Len Sparse Dense Speedup
128, 128 70.23 52.94 1.33x
128, 1024 69.11 52.00 1.33x
1024, 128 68.06 51.10 1.33x
1024, 1024 67.41 50.37 1.34x
L20
Inp Len, Out Len Sparse Dense Speedup
128, 128 54.75 29.86 1.83x
128, 1024 53.81 29.57 1.82x
1024, 128 52.49 29.18 1.80x
1024, 1024 51.64 28.94 1.78x
@@ 5 Conclusion
In this paper, we introduce the Adaptive Sparse Trainer (AST), a novel and efficient training pipeline for semi-structured sparse models. AST effectively narrows the precision gap between dense and sparse LLMs in terms of perplexity and accuracy on zero-shot tasks, while keeping training costs minimal. Our results demonstrate that pruning LLMs is feasible with minimal performance loss on knowledge-intensive tasks, and that large semi-structured sparse models can outperform dense models of similar decoding speed when supported by the appropriate software and hardware. Although our findings contribute to advancing the retraining of pruned models with billions of parameters, we acknowledge that the limited number of training tokens, due to computational constraints, remains an area for future exploration. Expanding this work to larger models or increasing the number of training tokens could provide valuable insights and further enhance the effectiveness of our methods.
Acknowledgements
The authors would like to thank Ziteng Wang, Bingrui Li and Pengle Zhang for helpful discussions and suggestions on code implementation. This work was supported by NSFC Project (No. 62376131).
@@ 6.1 Compressing Models with Different Semi-structured Sparsity Pattern
Through mathematical formulation (provided in Section 2 of the Appendix), we can provide an approximate upper bound for the compression ratio as well as a predicted actual compression ratio when extending to sparsity patterns with larger $N$ and $M$
@@ 6.4 SLoRB Initialization
Table 10: Wikitext perplexity for quantized sparse models.
LLAMA2 OPT GPT2
Method
Theoretical
Compression Rate
7B 125M 1.3B 124M 350M 774M 1.5B
Dense 1.0x 5.12 27.76 14.62 29.95 21.72 19.43 17.40
AWQ-4bit 0.125x 5.68 29.12 14.95 31.93 22.66 19.89 17.80
AWQ-2bit 0.0675x 2.2e5 216.11 53.42 751.15 453.45 70.36 46.17
AST-Naive-8bit
0.125x 6.37 30.26 15.86 32.36 23.66 21.29 18.34
AST-Naive-4bit
0.0675x 6.48 31.28 16.05 34.00 24.39 21.72 18.60
AST-Boosted-4bit
0.078x 6.25 29.88 15.63 32.76 23.87 21.40 18.31
@@ 6.5 Different Distillation Function
6.6 Adding Decay for AdamW Optimizer
Recent work (Hu et al. 2024) has shown that applying SR-STE with momentum-based optimizers can weaken the effectiveness of the decay term, resulting in suboptimal performance. In earlier implementations, SR-STE added the decay term directly to the weight. Specifically, we denote the gradient with respect to weight $W_{t}$ at iteration $t$ as $g_{t}$ which is further adjusted using the AdamW optimizer, followed by the weight update using:
$\text{AdamW}(g_{t})=\frac{(u_{t}\beta_{1}+(1-\beta_{1})g_{t})}{(1-\beta_{1}^{t})(\sqrt{v_{t}}+\epsilon)}$ (16)
$W_{t+1}\leftarrow W_{t}-\gamma_{t}\left(\text{AdamW}(g_{t})+\lambda_{W}(\overline{m(W_{t})}\odot W_{t})\right).$ (17)
where $u_{t}$ and $v_{t}$ are the first-order and second-order momenta of $g_{t}$, respectively. $\gamma_{t}$ is the current learning rate.
However, results indicate that this approach causes frequent mask oscillations details can be found in previous work (Hu et al. 2024). To mitigate this issue, the decay term is instead applied directly to the weight using the following update rule:
$\tilde{g_{t}}\leftarrow g_{t}+\lambda_{W}(\overline{m(W_{t})}\odot W_{t}),$ (18)
$W_{t+1}\leftarrow W_{t}-\gamma_{t}\text{AdamW}(\tilde{g_{t}}),$ (19)
We make further improvement to this method, since the decay term is only dependent to the parameter’s current value it should not be entangled with first order momentum. Therefore we employ the following update role:
$u_{t}=u_{t-1}\beta_{1}+(1-\beta_{1})g_{t-1}$ (20)
$\tilde{u_{t}}=u_{t}+\lambda_{W}(\overline{m(W_{t})}\odot W_{t})$ (21)
Furthermore we use the decayed gradient to calculate the second-order momentum and update weights accordingly.
$v_{t}=v_{t-1}\beta_{2}+(1-\beta_{2})\tilde{u_{t-1}}^{2}$ (22)
$W_{t+1}\leftarrow W_{t}-\gamma_{t}\frac{\tilde{u_{t}}}{(1-\beta_{1}^{t})(\sqrt{v_{t}}+\epsilon)}.$ (23)
@@ 6.8 Model Compression Using AWQ
In this section, we present the perplexity results for quantized models using AWQ. As shown in Table 10, our model maintains its performance even under extreme compression. For the AST-Boosted model, we also quantized the additional SLoRB parameters.
6.9 Frequently Asked Questions
This section examines and addresses key concerns raised in prior discussions.
Justification of Knowledge Distillation
One concern is that the use of distillation may introduce additional computational costs. However, we have found that, under the same total FLOPs, distillation methods still outperform non-distillation methods. We present a performance comparison under identical computational costs, and the results in Table 11 demonstrate that the distilled model significantly outperforms the non-distilled model.
Table 11: Perplexity for distillation and distillation-free methods under the same computational cost.
Module Tokens Total(TFLOPs) PPL
LLaMA2 w/ Distil 5B 3e8 5.97
LLaMA2 w/o Distil 7.5B 3e8 6.12
Justification for SLoRB
One concern is the timing of training the additional SLoRB component. Through experimentation, we observed that fine-tuning the LoRA component after initial training did not yield further performance improvements. We hypothesize that this is because the model has already converged on the pretraining dataset, making it difficult to further enhance performance through fine-tuning the LoRA module. In contrast, keeping both the model and LoRA modules unfrozen during retraining offers greater flexibility and potential for improvement. To support this hypothesis, we present the perplexity results for the GPT2 model in Table 12.”
Table 12: SLoRB ablation results on GPT2.
Model SLoRB w/o LoRA Finetune Aft. Retrain Dense
GPT2 31.13 32.23 32.07 29.95
Mask Selection
We present computational complexity and perplexity results for different mask selection criteria, such as the Wanda and SparseGPT metrics, in Table 13. The results demonstrate that the magnitude metric outperforms the others in both performance and speed.
Table 13: Wikitext perplexity for quantized sparse models.
Metric PPL Complexity
Magnitude 32.23 O($h^{2}$)
SparseGPT 32.58 O($h^{3}$)
Wanda 32.51 O($h^{2}$)
Effects of $\alpha$
We provide results using different methods on full GPT2 retraining and LLaMA2 retraining with limited training tokens in Table 14. When $\alpha=1$, meaning we only use KD loss, we observe slightly better performance on smaller models. However, it performs less effectively on larger models like LLaMA2-7B, where tuning $\alpha$ leads to improvements in performance. This suggests that $\alpha$ plays a more critical role in optimizing results for larger-scale models.
Table 14: Wikitext perplexity for different $\alpha$ on LLaMA2-7B and GPT2 model.
Model Tokens Trained $\alpha=0$ $\alpha=\frac{1}{3}$ $\alpha=\frac{2}{3}$ $\alpha=1$
LLaMA2-7B 3B 6.26 6.25 6.22 6.62
GPT2 5B 32.18 32.23 33.57 40.34